Grade 12Mechanics

a cube of ice floats partly in water and partly in oil .find the ratio of the volume of ice immersed in water to thatin oil. specific gravity of oil is 0.8 and that of ice is 0.9?
in this question how volume of ice become sum of the volume of oil and volume of water

apsara
8 Years agoGrade 12

1 Answer

Arun
8 Years ago

wieght of ice cube should be equal to the sum of wieght of oil replaced and  wight of water replaced by ice cube

let the volume in water is v1 and in oil is v2

0.9 * g * (v1+v2) =1* g*v1 + 0.8 * g* v2

or  0.9(v1/v2 +1) = v1/v2 +0.8

or v1/v2 =1